OpenStudy (anonymous):

solve 2sin^2x-2cos^2x-1=0

OpenStudy (anonymous):

2sin^2 (x) -2cos^2 (x) -1=0 so sin^2 (x) -cos^2 (x) =1/2 then sin^2(x)+cos^2(x)=1=2cos^2x+1/2 and 1/4=cos^2x or cosx = plus/minus 1/2

OpenStudy (anonymous):

You should be able to get the solutions from there.....

OpenStudy (anonymous):

Maybe of the form pi/4 +kpi/2 where k is an integer....